# Near-degeneracy effects in Quadrupolar Mixed Modes

### v1

@@ 4.2.1 KIC 7341231
In Figure 2, we show our global fit of the mixed modes of KIC 7341231, and highlight in panels (b) and (c) the identification of $\ell=2$ modes on two consecutive radial orders. We vet the asymptotic formula and fitting framework by comparing the results with the work of 2017A&A...605A..75D, which made use of complete evolutionary models of the star to analyse the radial order represented in panel (b). We show that our fit is very successful at catching all visible $\ell=2$ modes in the PSD, as first identified in Fig.1 of 2017A&A...605A..75D in the neighbourhood of $380$ µHz.
@@
In order to assess the gain from adding constraints from $\ell=2$ modes to the fitting procedure, we estimate the Bayesian Information Criterion (BIC, 10.1214/aos/1176344136) comparing the fit performed with $\ell=2$ mixed modes and $\ell=2$ p modes only. The inclusion of mixed $\ell=2$ modes in the model leads to a BIC that is $4,341$ points lower than when using only mixed $\ell=1$ modes and pressure $\ell=2$ modes on the target KIC 7341231. This substantial improvement in BIC indicates very strong statistical support for the extended model, suggesting that it provides a significantly better representation of the observed oscillation data. As a result, the physical inferences drawn from this model (such as internal rotation rates) can be considered significantly more robust and better constrained than those inferred from mixed $\ell=1$ modes only.
@@ 4.2.2 KIC 8179973
We then perform the first fit of the radial and $\ell=1,2$ mixed-mode pattern for KIC 8179973. This star presents a particularly asymmetric pattern in the $\ell=2$ regions, as the multiplets are very close to each other (see panels (a) and (b) of Fig. 3). The asymptotic prescription (Eq. 26) yields a very good interpretation of the PSD, with all visible modes being accurately represented (see also Fig. 10). The values of $\Delta\Pi_{1},\varepsilon_{g,1}$ and $q_{1}$ are also in the range of expected values from scaling relations and measurements of (2017A&A...600A...1M; 2023ApJ...954..152K). The small number of visible mixed $\ell=2$ modes is expected from the small mixed mode density parameter (2018A&A...616A..24G) $\mathcal{N}_{2}={\Delta\nu}/({\Delta\Pi_{2}\nu_{\text{max}}^{2}})\approx 3.4$ and the fact that the coupling factor $q_{2}$ is very small. Furthermore, for this star, we have used a much less informative prior for $q_{2}$ as in the fit of KIC 7341231, which shows the robustness of our fitting method. In addition, $\Delta\Pi_{2}$ and $\Delta\Pi_{1}$ are compatible within $2\sigma$ following Eq. (36). From the global fit, we obtain $\Omega_{\rm core}=677\pm 4$ µHz and $\Omega_{\rm env}=43\pm 6$ µHz, which are within the expected range for a young red giant star (2018A&A...616A..24G). Like for KIC 7341231, the uncertainties we obtain on KIC 8179973 are significantly smaller than those typically reported in similar studies (see, e.g., 2018A&A...616A..24G; 2014A&A...564A..27D). This improvement highlights the strength of our global fitting procedure, which (i) minimises the uncertainties by simultaneously exploiting information from multiple modes, and (ii) naturally mitigates the ambiguity between noise and genuine power excesses thanks to the intrinsic pattern constraints of the analysis. Compared to a fit that includes only the pure $\ell=2$ pressure modes, the BIC derived from the full asymptotic framework is lower by $3,300$ points. This provides strong statistical evidence that the prescription including mixed quadrupole modes and incorporating near-degeneracy effects, offers a superior inference of the underlying physical parameters. This measurement of near-degeneracy effects in a second early red giant paves the way for a fully automated fitting pipeline based on asymptotic mixed-mode parameters across $\ell$.
@@ 5 Conclusions
Our results for $\Delta\Pi_{1}$ and $\Delta\Pi_{2}$ in KIC 8179973 are consistent within $2\sigma$. However, in KIC 7341231, the values agree within a few tenths of seconds with non-overlapping errors, suggesting either a small deviation from the asymptotic theory or the influence of surface effects that were not fully accounted for in this work. In addition, our asymptotic formulation, including the $\ell=2$ modes, has enabled us to put stronger constraints on the rotation rates of our targets. Indeed, in KIC 7341231 we measured an envelope (resp. core) rotation rate of $57\pm 11$ nHz (resp. $777\pm 10$ nHz) with only $\ell=1$ modes, while with the addition of $\ell=2$ modes combined with a global fitting framework, we obtain an envelope (resp. core) rotation of $53\pm 6$ nHz (resp. $781\pm 4$ nHz). Those values are compatible with each other, and we also observe that the inclusion of the mixed $\ell=2$ modes makes the uncertainty drop by a factor of $\approx$2 ($2.5$ for the core rotation and $1.8$ for envelope rotation). This is mostly due to the inclusion of more modes in the fitting procedure. There might also be a different contribution from the sensitivity kernels of the $\ell=2$ modes, in particular in the p-mode cavity, as the $\ell=2$ modes have a shallower turning point, opening the way for future inversions. This strengthening of the constraints on rotation rates is further supported by the fact that including the mixed $\ell=2$ modes made the BIC of the fit of KIC 7341231 (resp. KIC 8179973) drop by more than $4,000$ (resp. $3,000$) points, which is a strong statistical evidence that the use of the new formulation allowing to fit $\ell=2$ mixed modes leads to better physical inference. In addition, as we have shown in Fig. 2, our fitting framework can distinguish between entwined split modes of different $\ell$, which shows the robustness of the fitting methodology.
@@
The asymptotic framework is also expected to simplify future studies of magnetic signatures and yield new results using not only the $\ell=1$ but also the $\ell=2$ modes to characterise magnetic fields in the RGB. We expect that the measurement of $q_{2}$ will help constrain the mode suppression phenomenon (e.g. 2016AAS...22740407F; 2017A&A...598A..62M; 2016PASA...33...11S) for $\ell=2$ modes as the loss in power of weakened non radial modes modes is expected to be linked to the transmission coefficient from the p mode cavity to the g mode cavity (2015Sci...350..423F; 2020MNRAS.493.5726L; 2025A&A...696A.134M). Low and radial-order-dependent values of $\Delta\Pi_{2}$ should be a hint of the presence of an internal magnetic field, as is it done with $\Delta\Pi_{1}$ (2020MNRAS.496.3829L; 2022A&A...667A..68B; 2023A&A...670L..16D). In a follow-up study (Liagre et al., in prep.), we will employ this new formulation and framework together with first-order magnetic effects (e.g. 2020MNRAS.496..620G; 2021A&A...650A..53B; 2021A&A...647A.122M; 2022A&A...667A..68B; 2022Natur.610...43L; 2023A&A...676L...9M; 2024ApJ...970...42B; 2024A&A...690A.217D; 2023A&A...680A..26L; 2024MNRAS.534.1060H; 2024ApJ...970...42B) to disentangle rotation-induced near-degeneracy asymmetries from magnetic ones in $\ell=2$ modes. This is particularly important, as it should provide stronger constraints on the magnetic field topology, as shown by 2024A&A...690A.217D.
@@
We anticipate that ongoing and upcoming space missions will continue to expand the sample of stars with good signal-to-noise ratios observed in this intermediate evolutionary phase. During its extended operations, the Transit Exoplanet Survey Satellite (TESS, 2014SPIE.9143E..20R) mission improved its temporal resolution, reducing the full-frame image cadence first to 10 minutes and later to 3 minutes. The TESS data are currently being calibrated for seismology (2024tkas.confE.123G, García et al., in prep.), which will enable large-scale and homogeneous seismic analyses. Moreover, the upcoming PLAnetary Transits and Oscillations of stars (PLATO; 2014ExA....38..249R) mission is expected to enable the detection of a large number of solar-like oscillators in the subgiant and early giant phases from the calibration sample, consistent with recent yield predictions (2024A&A...683A..78G). Using this global asymptotic description along with the Bayesian fitting framework, we expect not only to characterize with an increased precision the rotation profile of stars exhibiting mixed $\ell=2$ modes but also of those more evolved stars, exhibiting near-degenerate $\ell=1$ modes (for instance relying on Eq. (53), see also 2025arXiv250926319A).
@@ Acknowledgements.
L. Bugnet and L. Einramhof gratefully acknowledges support from the European Research Council (ERC) under the Horizon Europe programme (Calcifer; Starting Grant agreement N101165631). While partially funded by the European Union, views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The authors acknowledge the great support and feedback provided during the redaction of this article by Pr. Rafael García and Pr. Savita Mathur. We would also like to thank Dr. Emily Hatt for her insights on uncertainty estimates. The authors also thank the members of the Asteroseismology and Stellar Dynamics group of the Institute of Science and Technology Austria (ISTA) for very useful discussions: L. Barrault, S.B. Das, K. Smith. This paper includes data collected by the Kepler mission and obtained from the MAST data archive at the Space Telescope Science Institute (STScI). Funding for the Kepler mission is provided by the NASA Science Mission Directorate. STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555.
Software: AstroPy (astropy:2013; astropy:2018), Matplotlib (Hunter:2007), NumPy (harris2020array), SciPy (2020SciPy-NMeth), emcee (2013PASP..125..306F), celerite (celerite), slepc4py (DALCIN20111124; 10.1145/1089014.1089019), KADACS (2011MNRAS.414L...6G)
@@ Appendix A Perturbative treatment of the near degeneracy effects
In this appendix, we show how we can use perturbative methods to approximate the eigenvalues of the quadratic eigenvalue problem (Eq. 8). This approach provides closed-form solutions that are valid in the regime of weak near-degeneracy i.e. when the off-diagonal entries of $R$ are small compared to the diagonal one but not negligible (typically one order of magnitude less). In addition, we demonstrate that it offers a new way to determine both the core and envelope rotation rates from a single near-degeneracy–asymmetric splitting.
@@
$\omega^{2}a_{i}+\omega\sum_{j}R_{ij}a_{j}+\omega_{0,i}^{2}=0\,,\forall i.$ (46)
@@
$\omega^{2}a_{i}+\omega_{0,i}\sum_{j}R_{ij}a_{j}+\omega_{0,i}^{2}=0\,,\forall i.$ (47)
@@ C.2 Priors on KIC 7341231
A global analysis using PyA2Z gives $\Delta\nu$, $\nu_{\text{max}}$, and a guess value of $\Delta\Pi_{1}$ using the stretched period method. The priors on those parameters were then chosen according to Sec C.1. Refined values of $\Delta\Pi_{1}$ and $\Delta\nu$ are obtained by fitting the mixed $\ell=1$ modes according to the method described in Sec. 3.4.Because the angle of inclination is high, we had access to the $m=0$ components of each quintuplet. Hence, we decided to use the method described in 3.3.1 to constrain $q_{2}$. To use this method, we needed to access the following parameters and their associated uncertainties: $\Delta\Pi_{2}$, $\Delta\nu$, $\nu_{a},\nu_{b}$ and the position of the $\ell=2$ p mode that is given by $\nu_{0}-\delta\nu_{0,2}$. Where $\nu_{a}$ and $\nu_{b}$ are the positions of the $m=0$ components of each multiplet and $\nu_{0}$ is the frequency of the closest radial mode. $\Delta\Pi_{2}$ and its associated uncertainty was then obtained following Sec.3.2. In order to obtain $\delta\nu_{0,2}$, we fitted a bimodal Gaussian model to the averaged collapse of the echelle diagram. We obtained ${\delta\nu_{0,2}}/{\Delta\nu}=0.120\pm 0.002$. We finally peakbagged $\nu_{a},\nu_{b}$ and $\nu_{0}$ and obtained $\nu_{a}=380.382\pm 0.023$ $\nu_{b}=382.627\pm 0.01 µHz$, and $\nu_{0}=384.514\pm 0.011$ µHz. Using those values and their uncertainties, we can use a Monte-Carlo method to sample the distribution of $q_{2}$ according to equation (40). We obtained $q_{2}=0.041\pm 0.003$. We decided to use a uniform prior on $q_{2}$ within 3 sigma of the obtained value.

### v2

@@ 4.2.1 KIC 7341231
In Figure 2, we show our global fit of the mixed modes of KIC 7341231, and highlight in panels (b) and (c) the identification of $\ell=2$ modes on two consecutive radial orders. We vet the asymptotic formula and fitting framework by comparing the results with the work of 2017A&A...605A..75D, which made use of complete evolutionary models of the star to analyse the radial order represented in panel (b). We show that our fit is consistent with all visible $\ell=2$ modes in the PSD, as first identified in Fig.1 of 2017A&A...605A..75D in the neighbourhood of $380$ µHz.
@@
In order to assess the gain from adding constraints from $\ell=2$ modes to the fitting procedure, we estimate the Bayesian Information Criterion (BIC, 10.1214/aos/1176344136) comparing the fit performed with $\ell=2$ mixed modes and $\ell=2$ p modes only. For that, we fit the data again with split mixed dipole modes and pressure $\ell=2$ modes as usually done in the community and compare it against the full fit of the present work. The inclusion of mixed $\ell=2$ modes in the model leads to a BIC that is 3,726 points lower than when using only mixed $\ell=1$ modes and pressure $\ell=2$ modes on the target KIC 7341231. Because differences in BIC approximate:
$\Delta\mathrm{BIC}=-2\ln(\mathrm{BF})\,$, (48)
where BF is the bayes factor; such a large $\Delta\mathrm{BIC}$ corresponds to overwhelmingly strong statistical evidence in favour of the extended model. We do not report the absolute BIC values themselves, as they are not tied to any absolute scale and carry no stand-alone interpretation; only their differences have statistical meaning for model comparison. This substantial improvement in BIC therefore indicates very strong support for including the mixed $\ell=2$ modes, suggesting that the extended model provides a significantly better representation of the observed oscillation data. As a result, the physical inferences drawn from this model (such as internal rotation rates) can be considered significantly more robust and better constrained than those inferred from mixed $\ell=1$ modes only.
@@ 4.2.2 KIC 8179973
We then perform the first fit of the radial and $\ell=1,2$ mixed-mode pattern for KIC 8179973. This star presents a particularly asymmetric pattern in the $\ell=2$ regions, as the multiplets are very close to each other (see panels (a) and (b) of Fig. 3). The asymptotic prescription (Eq. 27) yields a very good interpretation of the PSD, with all visible modes being accurately represented (see also Fig. 11). The values of $\Delta\Pi_{1},\varepsilon_{g,1}$ and $q_{1}$ are also in the range of expected values from scaling relations and measurements of (2017A&A...600A...1M; 2023ApJ...954..152K). The small number of visible mixed $\ell=2$ modes is expected from the small mixed mode density parameter (2018A&A...616A..24G) $\mathcal{N}_{2}={\Delta\nu}/({\Delta\Pi_{2}\nu_{\text{max}}^{2}})\approx 3.4$ and the fact that the coupling factor $q_{2}$ is very small. Furthermore, for this star, we have used a much less informative prior for $q_{2}$ as in the fit of KIC 7341231, which shows the robustness of our fitting method. In addition, $\Delta\Pi_{2}$ and $\Delta\Pi_{1}$ are compatible within $2\sigma$ following Eq. (38). From the global fit, we obtain $\Omega_{\rm core}=671\pm 8$ µHz and $\Omega_{\rm env}=49\pm 8$ µHz, which are within the expected range for a young red giant star (2018A&A...616A..24G). Like for KIC 7341231, the uncertainties we obtain on KIC 8179973 are significantly smaller than those typically reported in similar studies (see, e.g., 2018A&A...616A..24G; 2014A&A...564A..27D). This improvement highlights the strength of our global fitting procedure, which (i) minimises the uncertainties by simultaneously exploiting information from multiple modes, and (ii) naturally mitigates the ambiguity between noise and genuine power excesses thanks to the intrinsic pattern constraints of the analysis. Compared to a fit that includes only the pure $\ell=2$ pressure modes, the BIC derived from the full asymptotic framework is lower by 3,185 points. This provides strong statistical evidence that the prescription including mixed quadrupole modes and incorporating near-degeneracy effects, offers a superior inference of the underlying physical parameters. This measurement of near-degeneracy effects in a second early red giant paves the way for a fully automated fitting pipeline based on asymptotic mixed-mode parameters across $\ell$.
@@ 5 Conclusions
Our results for $\Delta\Pi_{1}$ and $\Delta\Pi_{2}$ in KIC 8179973 are consistent within $2\sigma$. However, in KIC 7341231, the values agree within a few tenths of seconds with non-overlapping errors, suggesting either a small deviation from the asymptotic theory or the influence of surface effects that were not fully accounted for in this work. Indeed, because the fit of asymptotic mixed modes is very reliant on the position of the pure p-mode, a difference in the surface effects affecting the $\ell=1$ and $\ell=2$ notional p modes would suffice to explain a small difference in $\Delta\Pi$. In addition, our asymptotic formulation, including the $\ell=2$ modes, has enabled us to put stronger constraints on the rotation rates of our targets. Indeed, in KIC 7341231 we measured an envelope (resp. core) rotation rate of $57\pm 11$ nHz (resp. $777\pm 10$ nHz) with only $\ell=1$ modes, while with the addition of $\ell=2$ modes combined with a global fitting framework, we obtain an envelope (resp. core) rotation of $53\pm 6$ nHz (resp. $781\pm 4$ nHz). Those values are compatible with each other, and we also observe that the inclusion of the mixed $\ell=2$ modes makes the uncertainty drop by a factor of $\approx$2 ($2.5$ for the core rotation and $1.8$ for envelope rotation). This is mostly due to the inclusion of more modes in the fitting procedure. There might also be a different contribution from the sensitivity kernels of the $\ell=2$ modes. Specifically, their shallower turning points offer distinct constraints from those of $\ell=1$ modes in the p-mode cavity, suggesting a potential for improved radial resolution in future inversions. This strengthening of the constraints on rotation rates is further supported by the fact that including the mixed $\ell=2$ modes made the BIC of the fit of KIC 7341231 and KIC 8179973 drop by more than 3,000 points, which is a strong statistical evidence that the use of the new formulation allowing to fit $\ell=2$ mixed modes leads to better physical inference. In addition, as we have shown in Fig. 2, our fitting framework can distinguish between entwined split modes of different $\ell$, which shows the robustness of the fitting methodology.
@@
The asymptotic framework is also expected to simplify future studies of magnetic signatures and yield new results using not only the $\ell=1$ but also the $\ell=2$ modes to characterise magnetic fields in the RGB. We expect that the measurement of $q_{2}$ will help constrain the mode suppression phenomenon (e.g. 2015Sci...350..423F; 2016Natur.529..364S; 2017A&A...598A..62M) for $\ell=2$ modes (2016PASA...33...11S) as the loss in power of weakened non radial modes modes is expected to be linked to the transmission coefficient from the p mode cavity to the g mode cavity (2015Sci...350..423F; 2020MNRAS.493.5726L; 2025A&A...696A.134M). Low and radial-order-dependent values of $\Delta\Pi_{2}$ should be a hint of the presence of an internal magnetic field, as is it done with $\Delta\Pi_{1}$ (2020MNRAS.496.3829L; 2022A&A...667A..68B; 2023A&A...670L..16D). In a follow-up study (Liagre et al., in prep.), we will employ this new formulation and framework together with first-order magnetic effects (e.g. 2020MNRAS.496..620G; 2021A&A...650A..53B; 2021A&A...647A.122M; 2022A&A...667A..68B; 2022Natur.610...43L; 2023A&A...676L...9M; 2024ApJ...970...42B; 2024A&A...690A.217D; 2023A&A...680A..26L; 2024MNRAS.534.1060H; 2024ApJ...970...42B) to disentangle rotation-induced near-degeneracy asymmetries from magnetic ones in $\ell=2$ modes. This is particularly important, as it should provide stronger constraints on the magnetic field topology, as shown by 2024A&A...690A.217D.
@@
We anticipate that ongoing and upcoming space missions will continue to expand the sample of stars with good signal-to-noise ratios observed in this intermediate evolutionary phase. During its extended operations, the Transit Exoplanet Survey Satellite (TESS, 2014SPIE.9143E..20R) mission improved its temporal resolution, reducing the full-frame image cadence first to 10 minutes and later to 3 minutes. The TESS data are currently being calibrated for seismology (2024tkas.confE.123G, García et al., in prep.), which will enable large-scale and homogeneous seismic analyses. Moreover, the upcoming PLAnetary Transits and Oscillations of stars (PLATO; 2014ExA....38..249R) mission is expected to enable the detection of a large number of solar-like oscillators in the subgiant and early giant phases from the calibration sample, consistent with recent yield predictions (2024A&A...683A..78G). Using this global asymptotic description along with the Bayesian fitting framework, we expect not only to characterize with an increased precision the rotation profile of stars exhibiting mixed $\ell=2$ modes but also of those more evolved stars, exhibiting near-degenerate $\ell=1$ modes (using the full formalism or for instance relying on Eq. (56), see also 2025arXiv250926319A).
@@ Acknowledgements.
We thank the referee for their careful and constructive report, which has substantially enhanced both the quality and clarity of the manuscript. L. Bugnet and L. Einramhof gratefully acknowledge support from the European Research Council (ERC) under the Horizon Europe programme (Calcifer; Starting Grant agreement N101165631). While partially funded by the European Union, views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The authors acknowledge the great support and feedback provided during the redaction of this article by Pr. Rafael García and Pr. Savita Mathur. We would also like to thank Dr. Emily Hatt for her insights on uncertainty estimates. The authors also thank the members of the Asteroseismology and Stellar Dynamics group of the Institute of Science and Technology Austria (ISTA) for very useful discussions: L. Barrault, S.B. Das, K. Smith. This paper includes data collected by the Kepler mission and obtained from the MAST data archive at the Space Telescope Science Institute (STScI). Funding for the Kepler mission is provided by the NASA Science Mission Directorate. STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555.
Software: AstroPy (astropy:2013; astropy:2018), Matplotlib (Hunter:2007), NumPy (harris2020array), SciPy (2020SciPy-NMeth), emcee (2013PASP..125..306F), celerite (celerite), slepc4py (DALCIN20111124; 10.1145/1089014.1089019), KADACS (2011MNRAS.414L...6G), sloscillations (2019MNRAS.488..572K; 2023ApJ...954..152K)
@@ Appendix A Perturbative treatment of the near degeneracy effects
Although not used for our study, we provide in this appendix an explicit approximation to the solution of the eigenvalue problem, as it could be used as a simpler and faster implementation of the near degeneracy effects when they are not strong. In particular, we show how we can use perturbative methods to approximate the eigenvalues of the quadratic eigenvalue problem (Eq. 9). This approach provides closed-form solutions that are valid in the regime of weak near-degeneracy i.e. when the off-diagonal entries of $R$ are small compared to the diagonal one but not negligible (typically one order of magnitude less). In addition, we demonstrate that it offers a new way to determine both the core and envelope rotation rates from a single near-degeneracy–asymmetric splitting.
@@
$\omega^{2}a_{i}+\omega\sum_{j}R_{ij}a_{j}+\omega_{0,i}^{2}a_{i}=0\,,\forall i.$ (49)
@@
$\omega^{2}a_{i}+\omega_{0,i}\sum_{j}R_{ij}a_{j}+\omega_{0,i}^{2}a_{i}=0\,,\forall i.$ (50)
@@ Appendix B Vetting the asymptotic formulation
To propagate the observational uncertainties of the mixed-mode parameters into the inferred rotational components, we performed a Monte-Carlo analysis with N=10, [^0] N=10{,}000 realizations. For each realization, we randomly drew the parameters $(\nu_{a},\nu_{b},\Delta\Pi_{2},\zeta_{a},\zeta_{b})$ from independent Gaussian distributions centered on the value reported in 2017A&A...605A..75D and with their respective uncertainties. For every random draw, we then minimized the $\chi^{2}$ difference between the observed multiplet frequencies and the model frequencies computed following Sec. 2.3.2, treating $\Omega_{\mathrm{core}}$ and $\Omega_{\rm env}$ as free parameters. The optimization was carried out using the L-BFGS-B algorithm, with the bounds $\nu_{c}\in[0,1]\penalty 10000\ \mu{\rm Hz}$ and $\nu_{\rm env}\in[0,0.1]\penalty 10000\ \mu{\rm Hz}$. Only successful minimizations were retained in the final sample. The resulting posterior distributions of $\nu_{c}$ and $\nu_{\rm env}$ were summarized by their median values and their 16th–84th percentile ranges, which we report as our final uncertainties.
@@ C.2 Priors on KIC 7341231
A global analysis using PyA2Z gives $\Delta\nu$, $\nu_{\text{max}}$, and a guess value of $\Delta\Pi_{1}$ using the stretched period method. The priors on those parameters were then chosen according to Sec C.1. Refined values of $\Delta\Pi_{1}$ and $\Delta\nu$ are obtained by fitting the mixed $\ell=1$ modes according to the method described in Sec. 3.2.6.Because the angle of inclination is high, we had access to the $m=0$ components of each quintuplet. Hence, we decided to use the method described in 3.2.4 to constrain $q_{2}$. To use this method, we needed to access the following parameters and their associated uncertainties: $\Delta\Pi_{2}$, $\Delta\nu$, $\nu_{a},\nu_{b}$ and the position of the $\ell=2$ p mode that is given by $\nu_{0}-\delta\nu_{0,2}$. Where $\nu_{a}$ and $\nu_{b}$ are the positions of the $m=0$ components of each multiplet and $\nu_{0}$ is the frequency of the closest radial mode. $\Delta\Pi_{2}$ and its associated uncertainty was then obtained following Sec.3.2.2. In order to obtain $\delta\nu_{0,2}$, we fitted a bimodal Gaussian model to the averaged collapse of the echelle diagram. We obtained ${\delta\nu_{0,2}}/{\Delta\nu}=0.120\pm 0.002$. We finally peakbagged $\nu_{a},\nu_{b}$ and $\nu_{0}$ and obtained $\nu_{a}=380.382\pm 0.023$ $\nu_{b}=382.627\pm 0.01 µHz$, and $\nu_{0}=384.514\pm 0.011$ µHz. Using those values and their uncertainties, we can use a Monte-Carlo method to sample the distribution of $q_{2}$ according to equation (42). We obtained $q_{2}=0.041\pm 0.003$. We decided to use a uniform prior on $q_{2}$ within 3 sigma of the obtained value.
@@
C.3 Priors on KIC 8179973
A global analysis using PyA2Z gives $\Delta\nu$, $\nu_{\text{max}}$, and a guess value of $\Delta\Pi_{1}$ using the stretched period method. The priors on those parameters were then chosen according to Sec C.1. Refined values of $\Delta\Pi_{1}$, $\Delta\nu$ and a value of $q_{1}$ are obtained by fitting the mixed $\ell=1$ modes according to the method described in Sec. 3.2.6. In KIC 8179973, the angle of inclination is such that the m=0 modes of the $\ell=2$ quintuplet are either absent or buried in the noise of the star. Hence, we cannot use the same method as the on used for KIC 7341231 to constrain $q_{2}$. However, we can still use simple analytical arguments to put a rough upper bound on $q_{2}$ as a function of $q_{1}$.
$q_{1}$ in itself is obtained by fitting the mixed $\ell=1$ modes according to the method described in Sec. 3.2.6. We obtain $q_{1}=0.197\pm 0.001$ which yields that $q_{2}<0.107$ using the arguments presented in Sec.3.2.4.1. We use a uniform prior for the orientation of axis of rotation of the star with respect to the line of sight. This translates to a prior $\propto\sin(\theta_{\mathrm{inc}})$ for $\theta_{\mathrm{inc}}$ between $0^{\circ}$ and $90^{\circ}$.
C.4 Cornerplots and comparison data/model
We compared the frequencies obtained within our asymptotic Bayesian framework to the frequencies obtained by 2012ApJ...756...19D; 2017A&A...605A..75D using a peakbagging code. Both the asymptotic model and the peakbagged frequencies agree, particularly around the central frequency ($\nu_{\text{max}}$) as can be seen on figure 5. This congruence is expected, as the asymptotic approximation becomes less precise further from $\nu_{\text{max}}$ due to increasing sensitivity to surface effects. Crucially, the physically informed nature of our method allows it to account for small excesses of power that are often indistinguishable from noise in standard peakbagging procedures.
Figure 5: Echelle diagram comparing the oscillation frequencies derived from the Bayesian asymptotic model (circles, this work) with the peakbagged frequencies of KIC 7341231 (squares, 2012ApJ...756...19D; 2017A&A...605A..75D, referred to as Dh12 in the legend of the figure). The markers for the model frequencies are rendered semi-transparent when the separation from a peakbagged frequency exceeds 1 µHz.